OpenStudy (anonymous):

Ashe went from 160 pounds to 180 pounds once he started a weight-training program at his gym. What was Ashe's percent increase in weight? A. 8% B. 11.1% C. 12.5% D. 20%

OpenStudy (anonymous):

Divide his change by his original weight. His change = 180 - 160 = 20 Original weight = 160 20/160 = 0.125 or 12.5%

OpenStudy (anonymous):

What grade are you in?

OpenStudy (anonymous):

Thank you! and 7th

OpenStudy (anonymous):

Oh im in 8th grade but im home schooled aka Florida Virtual and im on a test right now

OpenStudy (anonymous):

Oh me too :)